OpenStudy (anonymous):

find the average rate of change of y=lnx on the interval [1/e, e] PLEASE HELP!

OpenStudy (turingtest):

average rate of change from x_1 to x_2=\[{y_2-y_1\over x_2-x_2}={\ln(e)-\ln(1/e)\over e-1/e}={1-(-1)\over e-1/e}={2e \over e^2-1}\]